An aircraft carrier is sailing at 15.4 m/s. The captain gives the order to make a 180 degree turn. If the carrier takes 180 seconds to complete the maneuver and follows a circular path while maintaining a constant speed, what is the radius of that circular path?

A 1800 m

B 440 m

C 220 m

D 880 m
Answer #1

X = V*t

pi *R = 15.4* 180

R = 15.4*180/pi = 882.355005 m D answer

Goodluck for exams Comment in case any doubt, will reply for sure..